If a sequence contains the terms tn=-2n^2-5n+10,what is the value of t25? I got 2385 can someone please explain?
Answer by stanbon(75887)

If a sequence contains the terms t(n)=-2n^2-5n+10,what is the value of t(25)? I got 2385 can someone please explain?
--------------------
t(25) = -2(25)^2-5(25)+10
= -2(625)-125+10
= -1250-125+10
= -1375+10
= -1365
-----------
Cheers,
Stan H.